one number is 7 more than a second number.Also,the sum of the square of the two numbers is 389.Find the two numbers.
Answer by mananth(16949)

let the number be x & x+7
..
x^2+(x+7)^2=389
..
x^2+x^2+14x+49=389
..
2x^2+14x+49-389=0
..
2x^2+14x-340=0
..
/2
x^2+7x-170=0
..
x^2+17x-7x-170=0
x(x+17)-7(x+17)=0
(x+17)(x-7)=0
x=-17 OR x=7
...
The numbers are -17,-10
7,14 does not satisfy the equation
...
m.ananth@hotmail.ca